OpenStudy (anonymous):

How would I solve this? 5X-12=2X+6

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

5 X-12 = 2 X+6 Add (12-2 X) to both sides: 3 X = 18 Divide both sides by 3: X = 6

OpenStudy (amistre64):

equations like to be balanced; so whatever we do to one side of the = sign has to be done to the other side if it going to change the value. 5x - 12 = 2x + 6 ; this is solved by getting x all alone; lets start +12 +12 with adding12. ---------------- 5x + 0 = 2x + 18 ; now we should subtract 2x for good measure -2x -2x ------------------ 3x + 0 = 0 + 18 ; and lets clean it up 3x = 18 ; now all we have to do to get that "x" all by itself /3 /3 is divide off that 3 ----------- x = 6